### v1

# Adaptive Attention Distillation for Robust Few-Shot Segmentation under Environmental Perturbations

Abstract
Few-shot segmentation (FSS) aims to rapidly learn novel class concepts from limited examples to segment specific targets in unseen images, and has been widely applied in areas such as medical diagnosis and industrial inspection. However, existing studies largely overlook the complex environmental factors encountered in real-world scenarios—such as illumination, background, and camera viewpoint—which can substantially increase the difficulty of test images. As a result, models trained under laboratory conditions often fall short of practical deployment requirements. To bridge this gap, in this paper, an environment-robust FSS setting is introduced that explicitly incorporates challenging test cases arising from complex environments—such as motion blur, small objects, and camouflaged targets—to enhance model’s robustness under realistic, dynamic conditions. An environment-robust FSS benchmark (ER-FSS) is established, covering eight datasets across multiple real-world scenarios. In addition, an Adaptive Attention Distillation (AAD) method is proposed, which repeatedly contrasts and distills key shared semantics between known (support) and unknown (query) images to derive class-specific attention for novel categories. This strengthens the model’s ability to focus on the correct targets in complex environments, thereby improving environmental robustness. Comparative experiments show that AAD improves mIoU by 3.3%–8.5% across all datasets and settings, demonstrating superior performance and strong generalization. The source code and dataset are available at: https://github.com/guoqianyu-alberta/Adaptive-Attention-Distillation-for-FSS.
I Introduction
Figure 1: The overview of this paper comprises: (a) a comparison of environmental difficulty between query and support images in Environment-Robust Few-Shot Segmentation; (b) the proposed Environment-Robust FSS Benchmark (ER-FSS), including its covered scenarios and datasets; (c) a comparison between the proposed Adaptive Attention Distillation (AAD) method (I) and existing approaches (pre-training segmentation methods (I) and traditional few-shot segmentation (FSS) methods (II)); and (d) comparative experimental results demonstrating the methodological advancement.
Image segmentation, a fundamental task in computer vision, aims to precisely delineate object boundaries and plays a critical role in domains such as medical imaging and aerospace applications [8, 4, 46, 6, 15]. However, building large-scale training datasets remains expensive and time-consuming, as it requires extensive manual annotation of segmentation masks. This challenge has motivated research into few-shot segmentation (FSS), which learns to segment new object categories from only a handful of labeled examples.
A common solution is the pretrain–finetune paradigm, where pre-trained segmentation models (e.g., Swin Transformer [27], SAM [20]) are adapted to new tasks using limited samples. Yet under severe data scarcity, fine-tuning often leads to overfitting. FSS addresses this limitation by learning transferable knowledge from a small set of support images to segment unseen query images. Most FSS methods adopt a meta-learning framework with Siamese or prototypical architectures [39, 44, 32, 49, 25, 2, 7, 13], enabling the model to extract class-level representations shared across instances. Recent studies [48, 39] demonstrate that such methods achieve nearly $70$ mIoU in the 1-shot setting on general benchmarks including PASCAL [37] and COCO [33].
Despite this progress, real-world conditions introduce complex environmental variations—such as illumination changes, cluttered backgrounds, object motion, and viewpoint shifts—that significantly increase the difficulty of query images compared to support images. These factors can obscure target boundaries, distort shapes, or cause severe blur, resulting in a sharp degradation of FSS performance outside controlled environments. Unfortunately, most existing studies, datasets, and models overlook these real-world challenges, ultimately limiting the practical deployment of FSS algorithms.
To address these challenges, the Environment-robust Few-shot Segmentation (ER-FSS) task (see Fig. 1) is introduced to improve the resilience of FSS models under environmental perturbations. The task targets typical hard cases in query images arising from complex real-world conditions, such as motion blur, small objects, camouflaged targets, and occlusion of key features. To better mirror practical usage, images exhibiting these challenges serve as query images, while simpler, cleaner samples captured in controlled settings are used as support images. Based on this setup, the ER-FSS benchmark is constructed, covering six scenario types and eight datasets. Unlike conventional datasets, ER-FSS more faithfully reflects model performance, generalization, and robustness across diverse environmental variations and domains, offering a realistic benchmark for evaluating both general and FSS models.
On the ER-FSS benchmark, evaluations of state-of-the-art pretrain–finetune and FSS models reveal that their robustness under environmental variation remains far below practical requirements. Further analysis shows that perturbations amplify feature discrepancies between same-class targets in query and support images, leading to attention drift—the model fails to focus on the correct category or its key visual cues. To mitigate this issue, Adaptive Attention Distillation (AAD) is proposed. The method repeatedly contrasts semantic information between support and query images and progressively distills class-specific attention through multiple refinement stages. By consolidating critical semantic cues, AAD enhances target localization and segmentation accuracy in complex environments. Experimental results demonstrate that AAD consistently outperforms existing pre-trained and FSS models, achieving an average IoU improvement of 3.3%–8.75% over current state-of-the-art (SOTA) approaches.
In summary, this work makes the following contributions:
  • Introduces the Environment-robust Few-shot Segmentation (ER-FSS) task and the accompanying ER-FSS benchmark, consisting of eight datasets across diverse scenarios to enable realistic, multi-scene evaluation of segmentation robustness.
  • Proposes Adaptive Attention Distillation (AAD), which iteratively contrasts semantic information between support and query images to distill class-specific attention, thereby improving target recognition and enhancing robustness under challenging environmental conditions.
  • Extensive experiments on ER-FSS show that AAD significantly outperforms existing FSS and pretrain–finetune models, achieving stronger generalization and higher robustness across a wide range of scenarios and settings.
II Related Work
II-A Few-Shot Segmentation (FSS)
FSS [37, 33, 24, 32, 2, 7, 42, 49, 14, 26, 31] is characterized by the absence of target domain data during training. The goal is to segment query images from an unseen domain using only a few annotated support images. OSLSM [37] was the first to tackle this problem, computing classifier weights for each query-support pair at evaluation. Inspired by ProtoNet [41], most modern FSS methods follow a meta-learning framework with dual branches: one extracts prototypes from support images, the other processes query images. These approaches generally fall into two categories: prototypical feature learning and relation-based methods.
Prototypical feature learning improves prototype representations to better separate foreground and background, enabling more accurate similarity measurement between support and query images. For example, PANet [44] employs prototype learning and a non-parametric decoder to create a consistent metric space. Relation-based methods, on the other hand, focus on improved measures of similarity after feature extraction.
However, most methods overlook a practical challenge: query images are often more complex than support images. When query images lack clear category cues—such as when key attributes are heavily occluded—FSS models may fail to identify the target or incorrectly segment unrelated objects. This paper extends FSS to a multi-domain English context, aiming to improve model robustness across diverse real-world settings. To address this, this paper introduces a new benchmark dataset and propose a novel approach tailored for robust multi-domain FSS.
II-B Related Datasets
To evaluate FSS model performance, most studies use PASCAL [37] and COCO [33], which cover common categories such as people, animals, vehicles, and indoor scenes, with COCO also including food, toys, and more. During testing, images for each novel class are randomly selected as support and query sets. However, these datasets do not adequately reflect the complexities of real-world applications.
To better assess FSS model generalization, the Cross-Domain Few-Shot Segmentation (CD-FSS) benchmark [21] was introduced, spanning the general domain (FSS-1000 [24]), medical (Chest X-ray [5]; ISIC [9]), and agricultural (Deepglobe [10]) domains. Despite this, CD-FSS remains limited in scope. Except for Deepglobe, FSS methods perform similarly on other datasets as in the general domain, indicating that CD-FSS does not fully capture the real challenges faced by current segmentation models. Like PASCAL and COCO, it also overlooks the impact of many difficult real-world cases on model generalization.
To address these gaps, we introduce the Environment-Robust Few-Shot Segmentation (ER-FSS) benchmark, designed to assess FSS under eight challenging scenarios, including camouflaged objects and small targets. ER-FSS provides a more comprehensive and realistic evaluation platform for FSS and broader segmentation models.
III Benchmark Dataset
Figure 2: The construction process of Environment-robust Few-shot Segmentation (ER-FSS) Benchmark: data collection phase (a) and manual annotation phase (b). Data statistics (c) and a visualization example with t-SNE for the evaluation benchmark datasets.
III-A Overview of the Benchmark Dataset
In this work, we introduce an Environment-Robust FSS Benchmark (ER-FSS) benchmark dataset, designed to serve as a comprehensive evaluation platform for FSS models and a wide range of segmentation algorithms under diverse and realistic environmental conditions. Compared with previous benchmarks, ER-FSS offers several key advantages: (1) it covers a broader array of specialized domains; (2) it features more meticulous data cleaning and manual annotation; and (3) it explicitly distinguishes between easy and difficult samples according to environmental complexity—challenging real-world samples are designated as “query“ (evaluation) images, while simple or laboratory-scenario images are categorized as “support“ (known) images.
Specifically, for (1), the ER-FSS benchmark includes eight evaluation datasets spanning six major domains: Animals (biology, segmenting 18 animal categories), Lunar Terrain (astronomy, segmenting surface elevations and depressions), Polyps (medicine, segmenting colon polyps in colonoscopy images), Eyeballs (medicine, segmenting retinal blood vessels), Road Cracks (industry, segmenting cracks on road surfaces), Steel Defects (industry, segmenting surface defects on steel structures), Leaf Diseases (agriculture, segmenting diseased regions on leaves), and City Satellite (geography, segmenting nine categories of objects, including buildings and roads, in satellite imagery).
For (2), every image in the dataset has undergone manual inspection for quality and label accuracy, accompanied by rigorous data cleaning to ensure the highest annotation standards. For (3), we identified common real-world challenges such as high similarity between the target and background in color or shape, targets appearing too small due to long-distance imaging, elongated target shapes, image blurring, and occlusion or absence of key target attributes. We summarize these difficulties into five challenging characteristics: camouflaged objects, small targets, elongated targets, missing attributes, and image blurring. Accordingly, every image was manually annotated as either an “easy sample“ or a “hard sample“ and assigned to the “support set“ or “query set“, respectively.
III-B Construction Process
As illustrated in Fig. 2, building the ER-FSS benchmark dataset comprises two primary stages: data collection and manual annotation.
Data Collection. In the data collection phase, we aimed to gather images from as many domains and sources as possible. To this end, we selected images from 17 diverse datasets. The sources for each dataset are as follows: Animals (MAS3K [23], DUTS [45], ECSSD [38], IS [22], COD10K [12]), Lunar terrain (ALLD), Polyp (CVC-ClinicDB), Eyeballs (DRIVE, STARE, STAREHRF [3], CHASE DB), Road cracks (CrackForest [40], CrackDataset [1]), Steel defects (MTD [19]), Leaf diseases (LDS), and City satellite (DeepGlobe 2018 [10], AIS). Following data collection, we performed rigorous cleaning and verification of the images. We ensured that each image had an accurate and valid corresponding mask label, removing any samples with incomplete or incorrect annotations.
Manual Annotation. We focus on two primary aspects in our annotation process: characteristics are defined as follows:
  • Manual Annotation of Class Labels: For images without category labels, we manually annotate the class texts and rectify errors in any existing labels.
  • Query/Support Sample Selection: Images exhibiting at least one of the challenging characteristics listed below are categorized as query images, while those that do not are defined as support images.
To ensure high-quality annotations, each image’s class label is verified by at least two annotators. The classification of simple versus difficult samples (i.e., support/query selection) is reviewed by at least three annotators, with repeated checks to minimize omissions and mislabeling.The challenging characteristics are defined as follows:
  • Camouflaged Objects: The target and background share similar visual attributes, such as color or texture, making it difficult for both models and humans to distinguish between them.
[…]
  • Elongated Targets: Targets with extremely elongated and irregular shapes (e.g., fine retinal blood vessels), which are difficult for models to accurately capture.
  • Missing Attributes: Crucial distinguishing features of the target are either occluded by other objects or missing due to incomplete capture.
  • Image Blurring: Reduced image clarity, often due to low resolution or motion blur, which makes target identification challenging.
Data Distribution and Statistics. To ensure the effectiveness of testing, the number of support images for all categories exceeds $20$, and the number of query images exceeds $10$. Moreover, to enhance dataset diversity, each dataset’s query images incorporate more than two challenging attributes. Additionally, we showcase visualization results of selected images from the Animal dataset after ViT [11] feature extraction with tSNE. As shown in Fig. 2, even within the same category—such as “snake,” “frog,” and “lizard”—the distribution of hard samples differs markedly from that of simple samples. Notably, hard samples of “frogs” are distributed more similarly to simple samples of “lizards” than to simple samples of their own class. This highlights a key challenge in real-world settings, where FSS and segmentation models often fail to identify targets correctly, or mistakenly classify them as background, resulting in segmentation errors. To address this, the ER-FSS benchmark proposed in this paper aims to offer a more realistic and application-oriented evaluation platform that better captures the complexities of real-world scenarios.
IV Method
Figure 3: The pipeline of the proposed Adaptive Attention Distillation (AAD) framework. AAD framework consists of four parts: the encoder, correlation leaner module, adaptive attention distillation learner, and the decoder.
IV-A Problem Setting
The environment-robust few-shot segmentation (ER-FSS) problem in this paper is formulated based on the classic FSS task. We assume a pre-training dataset $(X_{P},Y_{P})$ and a target domain dataset $(X_{T},Y_{T})$, where $X$ represents the input data distribution and $Y$ represents the corresponding segmentation labels. The model learns segmentation meta-knowledge from the extensive pre-training data $D_{P}\subseteq(X_{P},Y_{P})$ and is evaluated on the target data $D_{T}\subseteq(X_{T},Y_{T})$. Similar to FSS, in our ER-FSS setting, the input data distribution of the pre-training domain, $X_{P}$, differs from that of the target domain, $X_{T}$, and their label spaces are disjoint, i.e., $X_{P}\neq X_{T}$ and $Y_{P}\cap Y_{T}=\emptyset$.
The training set, denoted as $D_{\text{train}}$, is constructed from $(X_{P},Y_{P})$, while the testing set, $D_{\text{test}}$, is derived from $(X_{T},Y_{T})$. During the testing phase, we adhere to the episodic paradigm [43]. Specifically, for a given $N$-way $K$-shot learning task, $D_{\text{test}}$ comprises multiple episodes. Each episode is constructed with (1) a support set of simple samples, $S=\{({I}^{s}_{i},{M}^{s}_{i})\}_{i=1}^{N\times K}$, and (2) a query set of hard samples, $Q=\{({I}^{q}_{j},{M}^{q}_{j})\}_{j=1}^{n}$, where $I$ is an image, $M$ is the corresponding segmentation mask, and $n$ is the number of query images. Note that in our experiments, the training process for both the FSS baselines and our framework also follows the episodic paradigm. In contrast, when assessing pre-trained models based on transfer learning, the training phase adheres to their native single-branch, end-to-end approach.
IV-B Method Overview
Current methods in FSS aim to extract knowledge of a new class from a support image to then segment a query image. However, these methods are challenged by the significant domain gap that arises when the target’s environment is complex, causing large intra-class variance between the support and query sets. For example, as illustrated in Fig. 2, a “rabbit” in a support image may be clearly depicted, while in a natural query scene, its features can be obscured by camouflage or occlusion. This variance leads to biased and non-generalizable knowledge being learned from the support image, causing models to fail at accurately localizing and segmenting difficult targets in query images.
To overcome this limitation, we introduce Adaptive Attention Distillation (AAD), a strategy designed to learn more robust class representations. AAD iteratively compares the core features of support and query images to “distill” a stable, generalizable class-level attention. This distilled attention effectively guides the model to focus on the target while differentiating it from the background. The adaptive nature of this process enables generalization to diverse and unseen environments, thus enhancing the overall environmental robustness of the FSS model. The proposed AAD framework, depicted in Fig. 3, is composed of four modules: a shared encoder, a correlation learner, an Adaptive Attention Distillation learner, and a decoder.
IV-C Shared Feature Encoder
As illustrated in Fig. 3(a), the AAD framework begins with a feature encoder. This module, which can be implemented with a standard backbone architecture such as ResNet, VGG, or a Vision Transformer, adheres to a parameter-sharing paradigm. This design choice is crucial for few-shot learning, as it ensures that both the support image $I^{s}$ and the query image $I^{q}$ are projected into a common, consistent feature space.
During each training iteration, a support-query pair $(I^{s},I^{q})$ is sampled from the training dataset $D_{\text{train}}$. These images are processed by the shared backbone to produce their respective feature representations, $f^{s}$ and $f^{q}$. To facilitate robust matching across different levels of abstraction, we extract multi-scale feature maps. Specifically, for each image, we obtain features $f_{i}$ at three different scales $i\in\{\frac{1}{8},\frac{1}{16},\frac{1}{32}\}$, where the scale indicates the spatial resolution relative to the original input image size. Thus, $f^{s}_{i}$ and $f^{q}_{i}$ denote the feature maps from the $i$-th scale, each with a dimension of $\mathbb{R}^{H_{i}\times W_{i}\times d_{i}}$. Concurrently, the ground-truth segmentation mask of the support image, $Y^{s}$, is downsampled to match the spatial dimensions of the multi-scale feature maps. This results in a set of support masks, $M^{s}_{i}$, for each scale. These masks are instrumental as they provide explicit, location-specific information about the target category within the support feature space, which is essential for guiding the segmentation of the query image.
IV-D Correlation Learner
As shown in Fig. 3(b), the correlation learning module is designed to establish an initial correspondence between the support and query images. This module is composed of multiple Aggregation Mask Modules (AMMs), each operating at a specific feature scale. The primary function of the AMM is to generate a coarse segmentation mask for the query image by transferring knowledge from the support set. This is achieved by leveraging a cross-attention mechanism, which has proven effective at identifying feature similarities across different inputs. The process begins with the AMM at each scale $i$ taking the support features $f^{s}_{i}$ and query features $f^{q}_{i}$ as input. The cross-attention mechanism computes the similarity between them, effectively assigning higher weights to regions in the query features $f^{q}_{i}$ that closely correspond to the target features present in $f^{s}_{i}$. Subsequently, the module incorporates the corresponding support mask $M^{s}_{i}$ to filter and refine these weighted features, producing a coarse mask that approximates the target’s location in the query image.
These multi-scale coarse segmentation maps serve a dual purpose. First, they are forwarded to the decoder and progressively upsampled to contribute to the final, high-resolution segmentation result. Second, and more critically for our AAD framework, they provide the initial target localization information required by the subsequent Adaptive Attention Distillation learner, which will further refine this understanding.
Specifically, for a given scale $i$, the AMM first reshapes $f^{s}_{i}$ and $f^{q}_{i}$ from $\mathbb{R}^{H_{i}\times W_{i}\times d_{i}}$ to a flattened format of $\mathbb{R}^{(H_{i}\times W_{i})\times d_{i}}$ to prepare them for matrix multiplication. The module then computes an attention-guided query mask using scaled dot-product attention, as defined in the following equation:
$\text{Attention}(f^{q}_{i},f^{s}_{i},M^{s}_{i})=\text{softmax}\left(\frac{f^{q}_{i}(f^{s}_{i})^{T}}{\sqrt{d_{i}}}\right)M^{s}_{i}.$ (1)
In this formulation, the query features $f^{q}_{i}$ function as the ”Query” (Q), while the support features $f^{s}_{i}$ serve as the ”Key” (K). The dot product $f^{q}_{i}(f^{s}_{i})^{T}$ calculates a raw similarity matrix between every feature vector in the query and every feature vector in the support. By applying the softmax function, we obtain a set of attention weights where query features that are highly similar to support features receive larger values. This process effectively highlights regions in the query image that share semantic content with the support image.
The resulting attention map is then multiplied by the downsampled support mask $M^{s}_{i}$, which acts as the ”Value” (V). Since $M^{s}_{i}$ contains binary values indicating the target object’s location (1 for the target, 0 for the background), this final matrix multiplication effectively ”filters” the attention scores. It retains and aggregates the attention weights corresponding only to the target category, thereby producing a coarse probability map for the target’s location within the query image, which we denote as the coarse query mask $M^{q^{\prime}}_{i}$.
IV-E Adaptive Attention Distillation Learner
While the Correlation Learner provides an initial, coarse localization of the target, its reliance on direct feature matching makes it susceptible to failure in ”hard” cases where significant appearance gaps exist between the support and query images (e.g., due to camouflage, motion blur, or viewpoint changes). To overcome this, a more robust mechanism is needed to distill the essential, class-discriminative semantics of the target category, independent of low-level feature variations. For this purpose, we introduce the Adaptive Attention Distillation (AAD) learner, as in Fig. 3(c). This module is designed to generate a set of compact, highly informative ”class queries” that encapsulate the core characteristics of the target category. Instead of performing dense, pixel-to-pixel comparisons, the AAD learner abstracts class-level information, enabling it to focus on the target’s fundamental properties while ignoring distracting background clutter and appearance shifts. The module consists of multiple Adaptive Query Generators (AQGs), which leverage a small set of learnable parameters to interact with and distill information from the foreground features of both the support and query images.
The process begins by isolating the foreground features, which is critical for ensuring the learner focuses exclusively on the target object. Using $f^{s}_{i}$ and $f^{q}_{i}$ from the encoder and the masks $M^{s}_{i}$ and $M^{q^{\prime}}_{i}$ (from the support set and the previous stage, respectively), we compute the foreground features $F^{s}_{i}$ and $F^{q}_{i}$ via element-wise multiplication:
$F^{s}_{i}=f^{s}_{i}\otimes M^{s}_{i},\quad F^{q}_{i}=f^{q}_{i}\otimes M^{q^{\prime}}_{i}.$ (2)
This operation effectively masks out the background, minimizing its influence on the subsequent learning process.
Next, we initialize a small set of $N$ learnable class queries, denoted as $q\in\mathbb{R}^{N\times l}$, where $l$ is a user-defined hidden dimension. In our experiments, we found that a small number of queries (e.g., $N<100$) is sufficient, making this a computationally lightweight operation. These queries act as ”information collectors,” tasked with distilling the most salient class semantics.
The core of the AQG is a cross-attention mechanism, but its application here is novel. The class queries $q$ act as the ”Query” (Q), while the support foreground features $F^{s}_{i}$ serve as the ”Key” (K), and the query foreground features $F^{q}_{i}$ serve as the ”Value” (V). This formulation is designed to answer the question: ”Which parts of the query foreground $F^{q}_{i}$ are most relevant, given the class context provided by the support foreground $F^{s}_{i}$?” The updated queries are computed as follows:
$\tilde{q}_{i}=\text{MLP}\left(\text{LayerNorm}\left(\text{softmax}\left(\frac{q(F^{s}_{i})^{T}}{\sqrt{d_{i}}}\right)F^{q}_{i}+q\right)\right),$ (3)
where the output is passed through a LayerNorm and a simple MLP (consisting of two linear layers and a ReLU activation) to refine the representation and reduce its dimensionality back to the original $N\times l$.
It is crucial to note that this process is fundamentally different from dense feature matching. The lightweight queries $q$ do not learn a pixel-wise correspondence. Instead, they treat the entire set of support foreground features $F^{s}_{i}$ as a unified representation of the target class. This abstraction is key to the method’s robustness; it allows the queries to capture the essential what (the class identity) rather than the incidental where (the exact pixel locations), making the learned representation resilient to intra-class variations and noise in the coarse query mask $M^{q^{\prime}}_{i}$.
Furthermore, we leverage the hierarchical nature of features extracted by deep networks. Low-level features (from earlier layers, e.g., at scale 1/8) typically encode rich edge and texture details, while high-level features (from deeper layers, e.g., at scale 1/32) capture more abstract semantic information. To harness this, the query distillation process is performed iteratively across the different scales. The updated queries from one scale, $\tilde{q}_{i}$, become the input queries for the next scale. This allows the class queries to progressively refine their understanding, starting with general structural information and culminating in high-level semantic knowledge. The final output, $\hat{q}$, is a set of highly discriminative class queries that have distilled the core, environment-invariant essence of the target category, ready to guide the final segmentation decoder.
IV-F Decoder and Loss Function
The final stage of the AAD framework is to synthesize the information gathered by the preceding modules into a precise segmentation map. As illustrated in Fig. 3(d), this is achieved by the decoder, which integrates the coarse localization from the Correlation Learner with the high-level semantic knowledge from the AAD learner.
The inputs to this stage are the multi-scale coarse query masks $M^{q^{\prime}}_{i}$ and the final set of discriminative class queries, $\hat{q}$. To effectively combine these, the class queries $\hat{q}$ are used to refine the coarse masks. Specifically, we perform an element-wise multiplication between $\hat{q}$ and each coarse mask $M^{q^{\prime}}_{i}$. This operation re-weights the mask features, amplifying regions that are consistent with the distilled class semantics and suppressing irrelevant areas. The resulting refined masks are then concatenated with the original coarse masks to form a rich, fused representation, $R_{q}$, as shown in Equation 4:
[…]
where $\otimes$ denotes element-wise multiplication.
Finally, this fused representation $R_{q}$ is processed by the decoder module. The decoder is a simple yet effective architecture composed of several convolutional and upsampling layers that progressively merge the multi-scale features and restore the representation to the original image resolution, producing the final segmentation prediction. The entire network is trained end-to-end. We optimize the model parameters by minimizing the standard Binary Cross-Entropy (BCE) loss between the predicted segmentation map and the ground-truth query mask.
V Experiments
TABLE I: Comparison with SOTA methods on 1-shot, 5-shot, and 20-shot setting on Road cracks (Industrial), Steel defects (Industrial), and Leaf diseases (Agriculture). The numbers in Bold indicate the best performance.
Backbone Method Road crack Steel defect Leaf diseases
1-shot 5-shot 20-shot 1-shot 5-shot 20-shot 1-shot 5-shot 20-shot
ResNet-50 HSNet [ICCV2021] [32] 2.25 2.58 3.65 6.35 6.72 16.46 18.75 20.39 20.63
CyCTR [NIPS2021] [49] 0.88 0.89 0.83 7.98 9.24 3.01 13.05 13.56 11.37
PFENet  [TPAMI2022] [42] 0.44 0.41 0.39 9.06 9.06 9.14 11.60 11.06 11.13
DCAMA [ECCV2022] [39] 3.36 5.23 6.26 7.20 7.13 11.80 19.80 22.00 23.29
DIaM [CVPR2023] [16] 0.95 0.32 0.13 5.82 6.76 22.74 0.85 1.65 4.06
HDMNet [CVPR2023] [34] 1.20 1.28 2.34 6.43 6.01 14.29 17.82 18.70 16.18
PFENet++ [TAPMI2024] [30] 0.00 0.91 0.00 5.69 6.98 7.70 17.91 16.91 18.89
HMNet [NIPS2024] [47] 1.20 1.28 2.34 6.43 6.01 15.36 20.87 21.72 24.12
ABCDFSS [CVPR2024] [18] 5.42 5.95 8.03 7.90 12.14 12.34 21.94 20.65 26.78
NTRENet++ [TCSVT2025] [26] 0.54 3.39 3.32 4.41 12.58 32.48 6.96 18.07 23.47
AAD (Ours) 8.15 10.22 11.67 10.44 16.66 24.05 24.57 29.03 30.95
ResNet-101 CyCTR [NIPS2021] [49] 0.20 0.02 0.07 5.20 3.97 9.89 17.18 16.44 16.14
RepriNet [CVPR2021] [2] 1.31 4.13 1.47 6.70 5.54 3.04 12.40 11.53 9.86
DCAMA [ECCV2022] [39] 1.55 1.60 1.62 8.08 9.06 17.52 22.91 26.47 28.23
SCCAN [ICCV2023][48] 0.48 2.27 1.12 9.72 18.16 14.21 19.23 18.23 16.05
AAD (Ours) 8.41 10.65 12.18 15.36 24.68 27.97 25.13 30.31 32.31
Transformer SAM [arxiv2023] [20] 1.02 1.02 1.02 5.55 5.55 14.84 15.51 15.51 15.51
DCAMA [ECCV2022] [39] 11.67 11.62 12.23 12.25 15.28 30.42 27.73 29.44 30.46
AAD (Ours) 10.60 12.71 13.20 15.36 25.39 38.08 34.37 39.10 41.22
TABLE II: Comparison with SOTA methods on 1-shot, 5-shot, and 20-shot setting on biology dataset (Animal) and medical datasets (Polyp and Eyeballss). The numbers in Bold indicate the best performance.
Backbone Method Animal Eyeballs Polyp
1-shot 5-shot 20-shot 1-shot 5-shot 20-shot 1-shot 5-shot 20-shot
ResNet-50 HSNet [ICCV2021] [32] 45.78 48.91 49.14 9.57 9.61 9.60 15.42 14.91 14.52
CyCTR [NIPS2021] [49] 46.18 47.76 44.25 9.06 9.15 9.00 12.93 12.88 12.81
PFENet  [TPAMI2022] [42] 32.89 33.43 32.74 9.06 9.06 9.14 13.48 13.73 13.81
DCAMA [ECCV2022] [39] 48.80 56.53 59.06 9.51 9.55 9.60 14.77 14.01 14.32
DIaM [CVPR2023] [16] 7.68 15.44 19.46 9.51 8.17 9.75 2.77 6.20 13.73
HDMNet [CVPR2023] [34] 44.28 48.95 50.71 9.04 9.12 9.11 12.77 13.84 13.40
PFENet++ [TAPMI2024] [30] 42.22 44.29 44.64 9.17 9.18 9.15 13.79 13.47 13.76
HMNet [NIPS2024] [47] 40.36 49.03 48.21 8.98 9.10 8.98 11.75 13.61 13.32
ABCDFSS [CVPR2024] [18] 25.78 33.81 34.82 12.02 11.78 12.05 15.34 14.89 14.29
NTRENet++ [TCSVT2025] [26] 38.66 43.22 43.89 4.21 7.64 10.88 18.29 18.49 18.62
AAD (Ours) 52.85 56.29 58.79 13.04 13.71 13.85 21.93 23.91 23.89
ResNet-101 CyCTR [NIPS2021] [49] 52.20 52.36 53.49 9.12 9.03 8.99 13.40 12.68 13.06
RepriNet [CVPR2021] [2] 46.21 49.84 49.74 9.88 11.05 12.36 16.41 23.73 30.97
DCAMA [ECCV2022] [39] 55.51 59.35 59.06 10.20 10.75 11.21 21.71 25.13 31.85
SCCAN [ICCV2023][48] 47.93 57.05 57.61 9.01 8.90 9.12 12.64 13.47 12.99
AAD (Ours) 61.11 63.71 64.89 11.15 12.00 12.33 28.98 37.57 42.45
Transformer SAM [arxiv2023] [20] 16.98 16.98 19.01 8.93 8.93 8.93 15.83 15.83 15.83
DCAMA [ECCV2022] [39] 61.91 64.88 66.25 9.86 9.85 9.87 33.28 26.28 31.70
AAD (Ours) 63.26 65.85 65.54 11.47 12.46 12.72 32.85 51.79 59.68
TABLE III: Comparison with SOTA methods on 1-shot, 5-shot, and 20-shot setting on Lunar terrain (Astronomy), and City satellite (Geography) datasets. The ’Mean’ refers to the average results across all eight datasets in the FSHS benchmark. The numbers in Bold indicate the best performance.
Backbone Method Lunar terrain City Satellite Mean
1-shot 5-shot 20-shot 1-shot 5-shot 20-shot 1-shot 5-shot 20-shot
ResNet-50 HSNet [ICCV2021] [32] 3.84 4.01 4.15 9.61 10.08 10.25 14.98 14.65 16.05
CyCTR [NIPS2021] [49] 4.78 5.70 5.42 6.64 6.69 6.33 12.98 13.84 12.19
PFENet  [TPAMI2022] [42] 7.02 7.65 7.98 4.91 4.66 4.50 11.06 11.13 11.10
DCAMA [ECCV2022] [39] 5.28 6.86 8.47 8.06 6.58 3.93 14.60 15.99 17.09
DIaM [CVPR2023] [16] 3.36 6.56 9.74 5.19 7.51 10.03 4.52 6.55 11.21
HDMNet [CVPR2023] [34] 3.46 5.16 4.93 5.16 6.73 9.21 12.30 13.66 14.91
PFENet++ [TAPMI2024] [30] 5.74 4.24 7.89 7.73 10.35 7.72 11.36 11.81 12.19
HMNet [NIPS2024] [47] 4.13 5.50 9.39 8.42 9.41 10.17 14.09 12.85 14.71
ABCDFSS [CVPR2024] [18] 6.87 7.77 8.49 8.13 9.43 9.91 11.50 12.94 14.79
NTRENet++ [TCSVT2025] [26] 6.96 9.20 10.73 0.97 2.01 1.98 9.00 12.73 16.15
AAD (Ours) 13.04 14.57 16.16 9.91 10.78 11.38 19.24 21.90 23.84
ResNet-101 CyCTR [NIPS2021] [49] 4.94 5.64 6.28 5.40 5.06 4.69 13.46 13.15 14.08
RepriNet [CVPR2021] [2] 3.41 9.00 12.35 8.74 10.33 10.65 13.13 15.60 16.35
DCAMA [ECCV2022] [39] 4.91 7.62 8.23 9.48 10.00 11.25 16.79 20.03 21.12
SCCAN [ICCV2023][48] 6.34 7.53 6.78 6.17 9.00 5.17 13.94 16.83 15.38
AAD (Ours) 9.35 10.42 14.13 10.28 10.91 11.43 21.22 25.03 27.21
Transformer SAM [arxiv2023] [20] 4.45 4.45 4.45 8.68 8.68 8.68 9.62 9.62 11.03
DCAMA [ECCV2022] [39] 8.16 12.85 15.33 9.65 11.47 12.20 21.48 25.10 28.29
AAD (Ours) 9.86 17.13 21.21 10.00 12.51 14.20 23.47 29.62 33.23
V-A Experiment Setup
Datasets. We utilize the general datasets PASCAL [37], MSCOCO [33], and FSS-1000 [24] with SBD augmentation as pre-training data, then evaluate the trained models on the proposed ER-FSS benchmark datasets, as proposed in Sec.III. Note that for a fair comparison, we exclude classes that overlap between the pre-train datasets and the evaluation datasets.
Training and Testing Strategy. For pre-trained segmentation models (SAM [20]), we adhere to their inherent transfer learning training strategy by conducting end-to-end training on the pre-training dataset for the entire model. During the evaluation phase, we fine-tune the pre-trained model using support images, followed by generating segmentation predictions for query images. Regarding FSS models (HSNet [32], CyCTR [49], PFENet [42], DCAMA [39], DIaM [16], HDMNet [34], RepriNet [2], SCCAN [48],PFENet++ [30], HMNet [47], ABCDFSS [18], NTRENet++ [26] ), we adopt a meta-learning training strategy and subsequently evaluate the trained models using a meta-testing strategy. In each evaluation, we compute the average mean-IoU over $2$ runs [43], each with different random seeds. Additionally, each run comprises $1,000$ tasks for each dataset across the evaluation benchmark, maintaining consistency with the setting in  [24].
Evaluation Metric. We assess segmentation performance using the metric of Mean Intersection over Union (mIoU), a measure defined as the mean IoUs across all image classes. To compute the IoU for each category, we utilize the formula $IoU=\frac{TP}{TP+FP+FN}$, where $TP$, FPFP, and $FN$ represent the count of true positive, false positive, and false negative pixels in the predicted segmentation masks.
Implementation Details. For a fair comparison, we employ Swin-transformer [27], ResNet-50 [17], and ResNet-101 [17] as feature extraction networks, all of which are initialized with the weights pre-trained on ILSVRC [36] and kept frozen during the training process, following the previous works [24, 2]. The input dimensions for support and query images are set at $384\times 384$. ResNet-50 and ResNet-101 feature maps have channel dimensions of 256, 512, 1024, and 2048, while Swin-transformer feature maps have dimensions of 192, 384, 768, and 1536. For ResNet-50-based ADD, the number of learnable initialization queries is set to 15, and the interaction with $\frac{1}{8}$, $\frac{1}{16}$, and $\frac{1}{32}$ dimensions of support and query features. In the case of the ResNet-101-based model, the number of queries is set to 20, with feature interactions at $\frac{1}{32}$ dimension Meanwhile, for the Swin-transformer-based model, 15 queries are used, with feature interactions at $\frac{1}{8}$, $\frac{1}{16}$, and $\frac{1}{32}$ dimensions. The decoder is configured with $2$ convolutional layers, and between each module, bilinear interpolation is applied to upsample the feature maps by a factor of $2$, resulting in a total of $2$ upsampling functions. These networks were implemented using PyTorch, with AdamW [29] as the optimizer, a learning rate of 1e41e-4, and a weight decay of $0.05$. During training, the batch size is set to $120$, and the training process ran on $8$ NVIDIA A800-SXM4-80GB GPUs in parallel, with subsequent evaluation on one GPU.
V-B Comparison with SOTA Models
As in Tab. III, and III, extensive evaluations across the eight datasets of the ER-FSS benchmark demonstrate that our proposed AAD framework consistently and substantially outperforms existing SOTA FSS and pretrain-finetune models across all backbones and settings. When using a ResNet-50 backbone, AAD achieves a mean mIoU of 19.24%, 21.90%, and 23.84% in the 1-shot, 5-shot, and 2 0-shot settings, respectively. This represents a significant margin over the next best method, DCAMA, which scores 14.60%, 15.99%, and 17.09%. The performance gap is even more pronounced with a Transformer backbone, where AAD achieves a mean mIoU of 23.47% (1-shot), a 3.51% improvement over DCAMA.
This lead is particularly evident in challenging industrial and agricultural domains, where environmental perturbations are severe. On the Road Cracks dataset, our ResNet-50 based AAD achieve 8.15% in the 1-shot setting, while most competing methods, including PFENet and CyCTR, score below 1%. For Steel Defects, AAD (ResNet-101) reaches 15.36% (1-shot), a 5.64% improvement over the strong baseline SCCAN. In the biology and medical domains, AAD also establishes new SOTA results. With a Transformer backbone on the Polyp dataset, AAD achieves a remarkable 32.85% and 51.79% mIoU in 1-shot and 5-shot settings, surpassing the previous best (DCAMA) by 10.53% and 25.51%, respectively. Even in scenarios where other methods perform well, such as the Animal dataset, AAD with ResNet-101 leads with 61.11% mIoU (1-shot), improving upon DCAMA’s 55.51%.
The superior performance of AAD stems from its novel architecture designed to explicitly tackle the environmental variance between support and query images. Unlike traditional FSS methods that rely on direct, and often brittle, feature matching, AAD introduces an adaptive distillation process. By generating abstract class queries and iteratively refining them using foreground information from both support and query images, our model learns to distill the core, invariant semantics of a target category. This allows it to maintain focus on the object’s essential characteristics even in the presence of significant appearance shifts caused by camouflage, motion blur, or occlusion. The Correlation Learner provides a strong initial localization, while the Adaptive Attention Distillation learner purifies this understanding. This leads to a more robust and generalizable representation that excels in the complex, real-world scenarios presented by the ER-FSS benchmark, setting a new standard for robust segmentation.
We present visualization results comparing our method with SOTA approaches in Fig. 5. The qualitative comparisons highlight the limitations of existing methods when faced with the challenging scenarios in the ER-FSS benchmark. For instance, both HMNet and PFENet++ struggle significantly with elongated and small targets, often resulting in a complete failure to locate the target objects in the query image. In contrast, AAD demonstrates superior robustness and adaptability across all these difficult cases. It not only successfully transfers accurate category recognition capabilities to hard samples but also exhibits a remarkable ability to precisely delineate object boundaries. By distilling core, environment-invariant class semantics, AAD effectively ignores background clutter and appearance variations. This allows it to accurately identify and segment challenging targets—whether they are small, elongated, or camouflaged—producing segmentation masks that are both complete and clean, proving its effectiveness in real-world conditions.
Figure 4: Visualization of query maps at different stages with Swin-transformer backbone. After CL refers to the output after the correlation learning module.
Figure 5: Comparison of segmentation results between our method and SOTA methods on various evaluation datasets and under multiple difficult scenarios.
V-C Further Analysis
TABLE IV: Ablation Study of correlation learning (CL) module and class discriminative information learner (ADD) with Swin-transformer. The baseline refers to HSNet [32].
Method Steel defect Leaf disease Polyp
1-shot 5-shot 1-shot 5-shot 1-shot 5-shot
Baseline 7.72 7.56 20.31 20.94 13.48 13.13
+CL 14.29 21.24 27.73 29.44 22.32 26.28
ADD 15.36 25.39 34.37 39.10 32.85 51.79
(1) Ablation Study for Proposed Modules. We present experimental results for the proposed elements, the correlation learner (CL) module and ADD. Tab. IV indicates that incorporating CL improves performance by 6%, 17.8%, 14.0%, 18.2%, 19.4%, and 38.5% on three datasets compared to the baseline. The additional inclusion of ADD results in further improvements of 7.6%, 17.8%, 14.0%, 18.2%, 19.4%, and 38.5% over the baseline. It can be proved that ADD enhances the results of both the baseline and baseline+CL, with more pronounced effects as the number of shots increases. This is attributed to ADD’s ability to utilize more support images to generate more accurate class discriminative information.
Moreover, Fig. 4 illustrates the query feature maps after the Backbone, CL, and ADD, demonstrating that the interactive learning of the CL module helps the model disregard background features in certain queries. The addition of ADD significantly enhances the model’s focus on target class information within the query features.
TABLE V: Ablation study for different numbers of random queries with Swin-transformer.
Queries(#) Steel defect Leaf disease Polyp
1-shot 5-shot 1-shot 5-shot 1-shot 5-shot
0 7.72 7.56 20.31 20.94 13.48 13.13
5 14.88 19.38 33.53 38.53 36.35 49.31
15 15.36 25.39 34.37 39.10 32.85 51.79
30 14.96 21.40 33.04 37.49 31.80 43.39
50 13.83 20.89 33.13 35.64 36.51 51.26
100 14.82 14.23 32.36 42.55 31.15 32.99
(2) Ablation Study for Numbers of Random Queries. Tab. V demonstrates the impact of varying numbers of learnable queries when the backbone is the Swin-Transformer. From the table, it is observed that as the number of queries increases from $5$, there is minimal difference in the 1-shot results, oscillating within a range of approximately $\pm 2\%$. The 5-shot results show an initial increase followed by a decline. Considering the increased computational resources associated with a higher number of queries, we choose 15 as the base parameter for ADD model.
TABLE VI: Ablation study for different combinations methods of support features $f^{s}$, query features $f^{q}$, support masks $M^{s}$, and query aggregation masks $M^{q^{\prime}}$.
Combination Animal Steel defect City satellite
1-shot 5-shot 1-shot 5-shot 1-shot 5-shot
Baseline 51.80 55.92 7.72 7.56 9.61 10.08
Maskadd (Eq.6) 61.35 62.92 14.38 24.93 9.41 11.59
Concat (Eq.5) 61.47 63.52 14.37 20.94 9.34 11.29
Ours (Eq.23) 63.26 65.85 15.36 25.39 10.00 12.51
(3) Ablation Study for Feature Combination Methods. In the class discriminative information learner, learnable queries interact with support features $f^{s}$, query features $f^{q}$, support mask $M^{s}$, and query aggregation masks $M^{q^{\prime}}$ to learn category information. We employ a combined approach using Eq. 2 and Eq. 3. Tab. VI presents two alternative feature fusion methods. Maskadd (Eq. 5) draws inspiration from Mask2former, where features are first attended to and then combined with the mask. Concat (Eq. 6) involves concatenating features and masks before feeding them into ADG.The results in the table demonstrate the comprehensive superiority of our approach over the other two methods, indicating that directly identifying foreground features allows the model to learn more accurate category information.
$\tilde{q}^{m}_{n}=MLP(softmax(M^{q^{\prime}}_{i}+\frac{\tilde{q}^{m}_{n-1}(f^{s}_{i}\otimes M^{s}_{i})}{\sqrt{d_{i}}}f^{q}_{i})+\tilde{q}^{m}_{n-1}).$ (5)
$\tilde{q}^{m}_{n}=MLP(softmax((M^{s}_{i},M^{q^{\prime}}_{i})+$ (6)
$\frac{\tilde{q}^{m}_{n-1}(f^{s}_{i},f^{q}_{i})}{\sqrt{d_{i}}}(f^{s}_{i},f^{q}_{i}))+\tilde{q}^{m}_{n-1}),$
Where $(M^{s}{i},M^{q^{\prime}}{i})$ refer to the concatenation of two features.
TABLE VII: Comparison of experimental results with different backbones.
Backbone Animal Polyp
1-shot 5-shot 20-shot 1-shot 5-shot 20-shot
VGG-16 43.61 47.00 47.76 19.29 21.66 21.91
ResNet-50 52.85 56.29 58.79 21.93 23.91 23.89
ResNet-101 61.11 63.71 64.89 28.98 37.57 42.45
ViT 52.83 52.84 53.67 20.72 20.72 20.72
Swin 63.26 65.85 65.54 32.85 51.79 59.68
(4) Comparison Results for Multiple Backbones. Simultaneously, we showcase the results of our method on various backbones, all of which are loaded with pre-trained weights from ImageNet. ViT [11], on the other hand, is initialized with pre-training parameters from CLIP [35]. From Tab. VII, it can be observed that the results are optimal for ResNet-101 and Swin-Transformer, showing stability across multiple datasets and settings.
TABLE VIII: Comparison in comprehensive N-shot settings.
Method Polyp
1-shot 5-shot 10-shot 20-shot 30-shot 50-shot
FCN [28] 2.36 2.98 3.30 4.17 7.99 14.50
SAM [20] 15.83 15.83 15.83 15.83 15.83 15.83
DCAMA [39] 22.32 26.28 28.82 31.70 33.10 34.95
ADD 32.85 51.79 56.64 59.68 61.09 62.25
(5) Comparison on More Settings of Shots. Also, we provide comparison results between our method and SOTA approaches in various shot settings, as shown in Tab. VIII. It demonstrates that ADD outperforms the second-ranked DCAMA by 10.6%, 25.51%, 27.82%, 27.98%, 27.99%, and 27.30% in 1-shot, 5-shot, 10-shot, 20-shot, 30-shot, and 50-shot settings, respectively. This confirms that ADD effectively utilizes the given support images, and the performance steadily improves as the shot number increases. In contrast, FCN and SAM show relatively minor performance improvements as the shot number increases, indicating that transfer learning-based pre-trained segmentation models struggle to transfer knowledge from simple samples to hard ones effectively.
TABLE IX: Comparison of Results and spend time of different K-shot inference methods on 20-shot with Swin-transformer.
[…]
Vote 64.81 53.25 13.20 41.22 2.89
Average 65.54 59.68 12.14 41.02 2.51
(6) K-shot Inference. Additionally, when facing $K$-shot Inference, there are two commonly used methods: the voting method, which averages predictions for each support image and query separately, and the average method which averages the features of $K$ support images to generate a segmentation result for the query. Tab. IX displays the results of these two methods on ADD and compares their testing times in the 20-shot scenario. It indicates that the averaging method takes less time and outperforms the other by 6.4% on the Polyp. The differences between the two testing methods are minimal for the other three datasets. In summary, the averaging method provides a higher cost-effectiveness.
VI Conclusion
In this paper, we introduced the Environment-Robust Few-Shot Segmentation (ER-FSS) setting to address a critical gap in existing research: the poor performance of few-shot segmentation models in complex, real-world conditions. To facilitate research in this area, we established the ER-FSS benchmark, a comprehensive collection of datasets designed to evaluate model robustness against environmental challenges like motion blur, camouflage, and viewpoint shifts. We then proposed the Adaptive Attention Distillation (AAD) method, an innovative framework that enhances environmental robustness by learning to distill core, class-discriminative semantics. By iteratively contrasting support and query features to generate a set of compact class queries, AAD effectively guides the model to focus on the essential characteristics of the target category, proving resilient to significant intra-class appearance variations. Extensive experiments demonstrate that our AAD method significantly outperforms existing state-of-the-art FSS and pretrain-finetune approaches on the ER-FSS benchmark, establishing a new baseline for developing practical and robust segmentation models.